OpenStudy (anonymous):

Find the integral of (lny)/(squareroot y) and evaluate at 81 and 64

OpenStudy (anonymous):

seems like an integration by part problem to me.

OpenStudy (anonymous):

I tried that and got \[((2x ^{1/2}*\ln(x))-4x ^{1/2}\] but I'm not sure if that's right

OpenStudy (anonymous):

got an idea how to start such a problem? \[ \large \int \ln x \frac{1}{\sqrt{x}}dx= \int \ln x \cdot x^{-\frac{1}{2}}dx \] from here on use integration by parts.

OpenStudy (anonymous):

\[ \Large 2 \sqrt{x} \ln x - 4 \sqrt{x} = 2\sqrt{x}(\ln x -2 )\]

OpenStudy (anonymous):

so you are right yes.

OpenStudy (anonymous):

when I try inserting the values in I'm having a little trouble

OpenStudy (anonymous):

would you take the equation and insert 81 into it and minus that whole thing by the same equation but with 64 inserted?

OpenStudy (anonymous):

That's why I tried to simplify the expression, I haven't inserted the values yes though. It wont be a pretty number in the end anyway from what I can tell, because of the natural log.

OpenStudy (anonymous):

yes that's how it works @danielleg

OpenStudy (anonymous):

First insert the higher values of x and calculate, then minus the lower bound of the integral evaluated at said bound.

OpenStudy (anonymous):

I did that and eventually got 124.859

OpenStudy (anonymous):

but that isn't right

OpenStudy (anonymous):

\[ \large \approx 8.55 \]

OpenStudy (anonymous):

when I submitted that into the test it said it was wrong and gave it all in ln but when I put their answer into my calculator it came out to the same thing you got so maybe I'll send my teacher a message

OpenStudy (anonymous):

oh well I understand that, maybe they want the EXACT result, the exact result is of course in form of natural log.

OpenStudy (anonymous):

\[ \Large 18(\ln81-2)-16(\ln64-2) \approx 8.5579 \]

OpenStudy (anonymous):

you can simplify that expression even more, most of the time when you are dealing with integrals they don't request the exact result.

OpenStudy (anonymous):

oh ok thank you very much for your help

OpenStudy (anonymous):

welcome